A COMPLETE STRUCTURE FOR THE GAUCHE ROTAMER OF 1,1,2,2-TETRAFLUOROETHANE FROM MICROWAVE SPECTROSCOPY

\underline{NORMAN C. CRAIG}, Department of Chemistry, Oberlin College, Oberlin, OH 44074; BELÉN MATÉ, ANGELA HIGHT WALKER, and RICHARD D. SUENRAM, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899.

The complete molecular structure of the high-energy, gauche rotamer of 1,1,2,2-tetrafluoroethane, Freon F134, has been determined using Fourier transform microwave spectroscopy. A total of five different ^{13}C- and ^{2}H-substituted isotopomers and the normal species have been studied. Unlike the nonpolar anti rotamer, the gauche form has a permanent electric dipole moment, which permits a direct rotational analysis. The electric dipole moment was also determined using the Stark effect of two low-J transitions. The resulting dipole moment is μ_{c} = 8.186(7) x 10^{−30} C.m [2.454(2) Debye]. Two recent ab initio calculations, one using the hybrid Hartree-Fock, density functional method (ACM)^{a} and the other an MP2/6-31G^{∗∗} model^{b}, are in reasonable agreement with experimental findings.

a M. Muir and J. Baker, J. Mol. Phys. 89, 211 (1996). b S. Papasavva, K.H. Illinger, and J.E. Kenny, J. Phys. Chem. 100, 10100, (1996).